There are 15 balls in a rectangular box 5 are red, 6 are yellow and the rest is green. If 2 balls are selected at random without replacement one after the other, what is the probability that they are: a) both green (b) of the same color.
Answer by Alan3354(69443)

There are 15 balls in a rectangular box 5 are red, 6 are yellow and the rest is green. If 2 balls are selected at random without replacement one after the other, what is the probability that they are:
a) both green
There are 4 green balls, 15 balls total.
The 1st is 4/15, the 2nd is 3/14.
(4/15)*(3/14) = 2/35
(b) of the same color.
That can be:
2 red
(5/15)*(4/14) = 20/210
2 yellow
(6/15)*(5/14) = 30/210
2 green = 12/210
------------------
Add those up = 62/210 = 31/105